# Quantum Aspects of Natural Top Quark Condensation

### v2

@@ II.4 Removing Relative Time
A canonical normalization of the $H_{i}(X)$ kinetic term in any frame (which leads to a normalized Noether current, such as $iH^{\dagger}\overleftrightarrow{\partial}_{\!\!\!X}H$) dictates a Lorentz invariant normalization constraint on $\phi(r)$:
@@
$1=M_{0}^{3}\int d^{3}r\;|\phi(\vec{r})|^{2},\qquad\text{which implies,}\qquad\int ZM_{0}\;dr^{0}=1$ (42)
@@
This can be viewed as a “prenormalization” to define the free bilocal field and it’s Noether charges main2 , before turning on the interaction. The factor $Z$ is not absorbed into the fields after turning on the interaction since we don’t want the coupling constant, $g_{0}^{2}$, to be renormalized at the classical level. Indeed, $Z$ can be viewed as an operator, $Z\sim\delta(M_{0}r^{0})$. We remark that all field theories, even local ones, are implicitly defined with a “canonical” prenormalization that defines the semiclassical action after which we turn on the interaction. Then, at loop level we are compelled to renormalize the fields to preserve the canonicality, which in turn renormalizes the coupling constants.
We then turn on the interaction with coupling constant $g_{0}$ of eq.(II.2), and the action, becomes,
$\!\!\!\!\!\!\!\!\!\!S=\!\!\int d^{4}X\biggl{(}|D_{X}H(X)|^{2}+|H(X)|^{2}M_{0}^{3}\!\!\int\!\!d^{3}r\;\left(-|\partial_{\vec{r}}\phi(\vec{r})|^{2}\right)+|H(X)|^{2}\int dr^{0}d^{3}r\left(M_{0}^{4}\;{2g_{0}^{2}N_{c}}D_{F}(2r^{\mu})|\phi(\vec{r})\;|^{2}\right)\biggr{)}$ (43)
Note that there is no $Z$ factor in the interaction term, and $|\partial_{r^{\mu}}\phi|^{2}=|\partial_{r^{0}}\phi|^{2}-|\partial_{\vec{r}}\phi|^{2}\rightarrow-|\partial_{\vec{r}}\phi|^{2}$. We therefore integrate over $r^{0}$ in the interaction term :
@@
$V_{0}(2r)\rightarrow-\frac{1}{M^{2}_{0}}\delta^{3}(2\vec{r})=-\frac{1}{8M^{2}_{0}}\delta^{3}(\vec{r})\qquad D_{F}(2r)\rightarrow\frac{1}{M^{2}_{0}}\delta^{4}(2r^{\mu})=\frac{1}{JM^{2}_{0}}\delta^{4}(r^{\mu})$ (46)
@@
$=M_{0}^{3}\!\!\int\!\!d^{3}r\;\biggl{(}|\partial_{\vec{r}}\phi(r)|^{2}-{g_{0}^{2}N_{c}M_{0}}\frac{e^{-2M_{0}|\vec{r}|}}{8\pi|\vec{r}|}|\phi(\vec{r})|^{2}\biggr{)}.$ (48)
@@
As discussed in detail in main , the solution of the Schrödinger-Klein-Gordon SKG equation for $\phi(r)$ indeed extends to large distances, $\phi(r)\sim e^{-|\mu|r}/r$ where $|\mu|<\!\!M_{0}$ near critical coupling. This dilutes the value of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$. We find that the Yukawa coupling $g_{Y}\propto\phi(0)$ and $\lambda\propto g_{Y}^{4}\propto|\phi(0)|^{4}$. Moreover, inputting the known value of the Lagrangian mass of the BEH boson in the symmetric phase, which is $|\mu|^{2}=-(88)^{2}$ GeV2, we find that the scale $M_{0}$ is $M_{0}\approx 6$ TeV. The quartic coupling, $\lambda$, will be determined at loop level and we find remarkable agreement with the SM as discussed below. The degree of fine-tuning of the theory is also suppressed by $\phi(0)$ in a subtle way. Rather than the naive result one would expect from the NJL model, $\delta g_{0}^{2}/g_{c}^{2}\sim|\mu|^{2}/M_{0}^{2}\sim 10^{-4}$, we now obtain a linear relation: $\delta g_{0}^{2}/g_{c}^{2}\sim|\mu|/M_{0}\sim 1\%$. To obtain these results we need the precise relationship between $g_{Y}$ and $\phi(0)$ which follows from the induced Yukawa coupling of the bound state to free fermions.
@@ II.5 The Induced Bound State Yukawa Interaction
We see in eq.(II.2) that the Yukawa interaction of the bound state with the free scattering state fermions is now induced from the $\sqrt{N_{c}}$ term, $S^{\prime}$:
@@
We therefore see that the induced Yukawa coupling to the field $H(x)$ in the pointlike limit of the potential (which should be a reasonable low energy approximation) is:
$g_{Y}=g_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0)$ (54)
Here the behavior of $\phi(0)$ is a suppression of $g_{Y}$ that is a power-law, $\sim\sqrt{|\mu|/M_{0}}$ for small $\mu$, near the critical coupling. This power-law behavior is significantly different than the slow RG (logarithmc) evolution in the old NJL-based top condensation model, which is why we now obtain $M_{0}\sim 6$ TeV and reduced fine-tuning $\sim$ few $\%$ (rather than $M_{0}\sim 10^{15}$ GeV and extreme fine-tuning $\sim 10^{-26}$).
@@ III Quantum Fermion Loop Corrections to the Semiclassical Theory
$\!\!\!\!\!\!\!\!\!\!S^{\prime}=\!\int\!d^{4}x\biggl{(}\widetilde{Z}DH^{\dagger}DH+M^{2}H^{\dagger}H-\frac{\lambda}{2}(H^{\dagger}H)^{2}\biggr{)}.$ (56)
@@
Hence, in the present analysis we will use the NJL limit in obtaining approximate bilocal loop integral results, but judiciously replacing $\delta^{4}(x-y)\rightarrow M_{0}^{2}D_{F}(x-y)$. Physically, the bilocal theory has a true momentum space cutoff of loop integrals $\sim M_{0}^{2}$, due to $D_{F}(q^{2})$, while in the NJL theory this cut-off is implemented in loop integrals “by hand.” The replacement of the $\delta^{4}(x-y)$ by $M_{0}^{2}D_{F}(x-y)$ approximately implements the true physics of the bilocal theory, allowing us to obtain a potential that, in turn, allows the “back-reaction” of the solution. We are invoking the “indistinguishability” of the $\delta^{4}(x-y)$ from the bilocal $M_{0}^{2}D_{F}(x-y)$, but maintaining the bilocality of the present theory. By extending the auxiliary field equation to include the additional terms of $W(\Omega)$, then integrating out $B_{i}$, we obtain the quantum loop corrections to the semiclassical action. These are the analogues of the terms in eqs.(56,57) generated by eq.(55), but now for the bilocal fields.
@@
$\int_{xy}\biggl{(}g_{0}\overline{\omega}_{R}(y)D_{F}(x-y)\omega_{iL}(x)B^{i\dagger}(x,y)+h.c.\biggr{)}$ (58)
@@
$+\!\int_{x...z^{\prime}}\!\!\!\![B^{\dagger}(x,y)B(x^{\prime},y^{\prime})]{\cal{G}}(x,y,x^{\prime},y^{\prime},w,z,w^{\prime},z^{\prime})[B^{\dagger}(w,z)B(w^{\prime},z^{\prime})]+$..., (59)
@@ III.1 Main Loop
$-\frac{1}{2}g_{0}^{2}N_{c}\!\int_{x...y^{\prime}}\!\!\![B^{\dagger}(x,y)D_{F}(x-y)\mathop{\rm Tr}(S_{F}(x-x^{\prime})S_{F}(y^{\prime}-y))D_{F}(x^{\prime}-y^{\prime})B(x^{\prime},y^{\prime})]$ (60)
@@
$S_{F}(x)=\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{i\ell\!\!\!/\!\,\,}{\ell^{2}+i\epsilon}e^{i\ell\cdot x}\qquad\makebox{and}\qquad\mathop{\rm Tr}\biggl{(}\frac{(1-\gamma^{5})}{2}\ell\!\!\!/\!\,\,\;\ell\!\!\!/\!\,\,^{\prime}\biggr{)}=2\ell\cdot\ell^{\prime}$ (61)
@@
$D_{F}(x-y)=-\int\frac{d^{4}q}{(2\pi)^{4}}\frac{e^{iq(x-y)}}{q^{2}-M_{0}^{2}}\rightarrow\int\frac{d^{4}q}{(2\pi)^{4}}\frac{e^{iq(x-y)}}{M_{0}^{2}}=\frac{1}{M_{0}^{2}}\delta^{4}(x-y)$ (62)
@@
Hence we have,
@@
$\!\!\!\!\!\!\!\!\!\!=-\frac{g_{0}^{2}N_{c}}{M_{0}^{4}}\!\int_{XX^{\prime}}\![\widetilde{B}^{\dagger}(X,0)\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{d^{4}\ell^{\prime}}{(2\pi)^{4}}\frac{2\ell\cdot\ell^{\prime}}{\ell^{2}\ell^{\prime 2}}e^{i(\ell-\ell^{\prime})(X-X^{\prime})}\widetilde{B}(X^{\prime},0)]$ (66)
@@
$=\frac{ig_{0}^{2}N_{c}}{8\pi^{2}M_{0}^{4}}\!\int_{\overline{X}}[\widetilde{B}_{0}^{\dagger}\widetilde{B}_{0}]\biggl{(}M_{0}^{2}+P^{2}\ln\left(\frac{M_{0}}{m}\right)\biggr{)}$ (67)
@@
We have obtained the analogues of the $M^{2}$ and $P^{2}$ terms in eqs.(56,57) acting on the auxiliary field. The $P^{2}$ term corresponds to a loop-generated kinetic term for the auxiliary field $B_{i}$, (see eq.(86)). In the framework of the NJL model and original top condensation theory, BHL , this corresponds the formation of a physical state, i.e., a resonance. We are interested presently in small $P^{2}$ so we will treat the $P^{2}\approx 0$ limit of the loop and return to this issue below.
@@
$W_{\cal{F}}(\Omega)=\frac{g_{0}^{2}N_{c}}{8\pi^{2}M_{0}^{2}}\!\int d^{4}X\;[\widetilde{B}^{\dagger}(X,0)\widetilde{B}(X,0)]\rightarrow\frac{g_{0}^{2}N_{c}J}{8\pi^{2}M_{0}^{2}}\!\int d^{4}Xd^{4}r\;\widetilde{B}^{\dagger}(X,r)\delta^{4}(2r)\widetilde{B}(X,r)$ (68)
@@
$W_{\cal{F}}(\Omega)=\!\int_{x...y^{\prime}}\!\!\!B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})\approx\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\int d^{4}x\;d^{4}y\;B^{\dagger}(x,y)D_{F}(x-y)B(x,y)$ (69)

### v3

@@ II.4 Removing Relative Time
A canonical normalization of the $H_{i}(X)$ kinetic term in any frame (which leads to a normalized Noether current, such as $iH^{\dagger}\overleftrightarrow{\partial}_{\!\!\!X}H$) dictates a Lorentz invariant normalization constraint:
@@
$1=M_{0}^{3}\int d^{3}r\;|\phi(\vec{r})|^{2},\qquad\text{which implies,}\qquad\int ZM_{0}\;dr^{0}=1,\qquad\text{(see below).}$ (43)
@@
This can be viewed as a “prenormalization” to define the free bilocal field and it’s Noether charges main2 , before turning on the interaction. The factor $Z$ is not absorbed into the fields after turning on the interaction since we don’t want the coupling constant, $g_{0}^{2}$, to be renormalized at the classical level. Indeed, $Z$ can be viewed as an operator, $Z\sim\delta(M_{0}r^{0})$. We remark that all field theories, even local ones, are implicitly defined with a “canonical” prenormalization that defines the semiclassical action after which we turn on the interaction. Then, at loop level, we are compelled to renormalize the fields to preserve the canonicality which, in turn, renormalizes the coupling constants.
We see that $S_{KTH}$ becomes,
$S_{KTH}\rightarrow\!\!\int\!\!d^{4}X\biggl{(}|D_{X}H(X)|^{2}+|H(X)|^{2}M_{0}^{3}\!\!\int\!\!d^{3}r\;\left(-|\partial_{\vec{r}}\phi(\vec{r})|^{2}\right)\biggr{)}.$ (44)
We then turn on the interaction with coupling constant $g_{0}$ of eq.(II.2), and the semiclassical action becomes,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{0}(H)=\!\!\int\!\!d^{4}X\biggl{(}|D_{X}H(X)|^{2}+|H(X)|^{2}M_{0}^{3}\!\!\int\!\!d^{3}r\;\left(-|\partial_{\vec{r}}\phi(\vec{r})|^{2}\right)+|H(X)|^{2}\!\!\int\!\!dr^{0}d^{3}r\left(M_{0}^{4}\;{2g_{0}^{2}N_{c}}D_{F}(2r^{\mu})|\phi(\vec{r})\;|^{2}\right)\!\biggr{)}.$ (45)
Note that there is no $Z$ factor in the interaction term, and $|\partial_{r^{\mu}}\phi|^{2}=|\partial_{r^{0}}\phi|^{2}-|\partial_{\vec{r}}\phi|^{2}\rightarrow-|\partial_{\vec{r}}\phi|^{2}$. If we were to define $z\equiv\int ZMdr^{0}$ , then we can rescale $\phi\rightarrow\phi/\sqrt{z}$ to have the compact normalization of eq.(43), but then the coupling constant becomes $g_{0}^{2}/z$. However, we want the theory to have at the classical level the same coupling $g_{0}^{2}$ as defined by the coloron exchange, e.g., we want to match the low energy scattering matrix elements that we would obtain in the conventional topcolor field theory to the bilocal theory. Hence $z=1$ is a matching condition of the underlying topcolor interaction to the bilocal theory.
We therefore integrate over $r^{0}$ in the interaction term :
@@
$V_{0}(2r)\rightarrow-\frac{1}{M^{2}_{0}}\delta^{3}(2\vec{r})=-\frac{1}{8M^{2}_{0}}\delta^{3}(\vec{r})\qquad D_{F}(2r)\rightarrow\frac{1}{M^{2}_{0}}\delta^{4}(2r^{\mu})=\frac{1}{JM^{2}_{0}}\delta^{4}(r^{\mu}).$ (48)
@@
The manipulation leading to eq.(47) may seem to be inconsistent, but it is a short-cut for a more formal limiting procedure. In a manifestly Lorentz invariant form of the internal wave-function action of eq.(45),
$\!\!\int\!\!d^{4}r\;\biggl{(}Z|\partial_{r^{\mu}}\phi(r^{\mu})|^{2}+\;{2g_{0}^{2}N_{c}}D_{F}(2r^{\mu})|\phi(r^{\mu})\;|^{2}\!\biggr{)},$ (49)
we want to take $\phi(r^{\mu})\rightarrow\phi(\vec{r})$, which is independent of $r^{0}$. However this appears to be inconsistent with the interaction term with the factor $D(2r^{\mu})$ which has nontrivial dependence upon $r^{0}$. The problem arises from having integrated out non-static colorons to produce $D_{F}(2r^{\mu})$. However to obtain a Hamiltonian (such as in functional Schrödinger picture) we go to static field configurations. To see this, consider a toy constituent, complex scalar, field theory:
$S=\!\!\int\!\!d^{4}x\biggl{(}(|\partial\varphi_{1}(x)|^{2}+|\partial\varphi_{2}(x)|^{2}-gM_{0}(\varphi_{1}(x)\varphi_{2}(x)A(r)+h.c.)+|\partial A(x)|^{2}-M_{0}^{2}|A(x)|^{.}\biggr{)}.$ (50)
The Hamiltonian for this theory takes the form
$H=\!\!\int\!\!d^{3}x\;\biggl{(}|\pi_{1}(\vec{x})|^{2}+|\pi_{2}(\vec{x})|^{2}+|\nabla\varphi_{1}(\vec{x})|^{2}+|\nabla\varphi_{2}(\vec{x})|^{2}+gM_{0}(\varphi_{1}(\vec{x})\varphi_{2}(\vec{x})A(\vec{x})+h.c.)$
$\qquad\qquad+|\pi_{A}(\vec{x})|^{2}+|\nabla A(\vec{x})|^{2}+M_{0}^{2}|A(\vec{x})|^{2}\biggr{)},$ (51)
where the canonical momenta are $\pi_{i}=\partial_{0}\varphi_{i}$ and $\pi_{A}=\partial_{0}A$. We are interested in the static limit of this theory so we impose the constraints, $\pi_{i}=\pi_{A}=0$, and there is then no Poisson bracket (hence no Dirac bracket). The static equation for $A(\vec{x})$ is, $\nabla^{2}A-M_{0}^{2}A=gM_{0}\varphi_{1}\varphi_{2}$, hence we can integrate out $A$, and write,
$H=\!\!\int\!\!d^{3}xd^{3}y\;\biggl{(}ZM_{0}|\nabla_{x}\varphi_{1}(x)|^{2}|\varphi_{2}(y)|^{2}+ZM_{0}|\nabla_{y}\varphi_{2}(y)|^{2}|\varphi_{1}(x)|^{2}-g^{2}M^{2}_{0}|\varphi_{1}(x)\phi_{2}(y)|^{2}\Delta(x-y)\biggr{)},$ (52)
where $1=ZM_{0}\int d^{3}y|\varphi_{i}(y)|^{2}$, and $\Delta(x-y)=V_{0}(\vec{x}-\vec{y})/2$. Now, replace $\varphi_{1}(x)\varphi_{2}(y)\rightarrow M_{0}\Phi(x,y)$ to obtain,
$H=\!\!\int\!\!M_{0}^{3}d^{3}xd^{3}y\;\biggl{(}Z|\nabla_{x}\Phi(x,y)|^{2}+Z|\nabla_{y}\Phi(x,y)|^{2}-\frac{1}{2}g^{2}M_{0}|\Phi(x,y)|^{2}V_{0}(x-y)\biggr{)}.$ (53)
We go to the barycentric coordinates, and factorize $\Phi(x,y)=\sqrt{2/J}\;\chi(X)\phi(\vec{r})$, [^7] A c-number product $\varphi_{1}(\vec{x})\varphi_{2}(\vec{y})$ cannot generally be rewritten in the factorized form, but quantum fields can lead to states that are so factorized, $\langle 0|\varphi_{1}(\vec{x})\varphi_{2}(\vec{y})|S\rangle\sim\chi(X)\phi(\vec{r})$.
$H=\!\!\int\!\!M_{0}^{3}d^{3}Xd^{3}r\;\biggl{(}Z|\nabla_{X}\chi(\vec{X})|^{2}|\phi(\vec{r})|^{2}+Z|\chi(\vec{X})|^{2}|\nabla_{\vec{r}}\phi(\vec{r})|^{2}-g_{0}^{2}N_{c}M_{0}|\chi(\vec{X})|^{2}|\phi(\vec{r})|^{2}V_{0}(2\vec{r})\biggr{)},$ (54)
with $g^{2}=g_{0}^{2}N_{c}$. We can do the same construction with the full coloron theory. The above example displays the origin of $Z$ and shows that there is no inconsistency in the final result of eq.(47) obtained by the short-cut which yields the single particle exchange interaction as it would be generated in the static frame. This is the correct procedure leading to a Hamiltonian for the functional Schrödinger equation of the field theory with the bound state (see the discussion in CJT of Kuti’s method). To recover the action formalism we would do a Legendre transformation from $H$ back to the Lagrangian $L$, restoring the time derivative on $\chi(X)$ and integrating $L$ over $X^{0}$.
@@
${\cal{M}}=M_{0}^{3}\!\!\int\!\!d^{3}r\;\biggl{(}|\partial_{\vec{r}}\phi(r)|^{2}-{g_{0}^{2}N_{c}M_{0}}\frac{e^{-2M_{0}|\vec{r}|}}{8\pi|\vec{r}|}|\phi(\vec{r})|^{2}\biggr{)}.$ (56)
@@
As discussed in detail in main , the solution of the Schrödinger-Klein-Gordon SKG equation for $\phi(r)$ indeed extends to large distances, $\phi(r)\sim e^{-|\mu|r}/r$ where $|\mu|<\!\!M_{0}$ near critical coupling. This dilutes the value of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$. We find that the Yukawa coupling $g_{Y}\propto\phi(0)$ and $\lambda\propto g_{Y}^{4}\propto|\phi(0)|^{4}$. Moreover, inputting the known value of the Lagrangian mass of the BEH boson in the symmetric phase, $|\mu|^{2}=-(88)^{2}$ GeV2, we find the composite scale (coloron mass) $M_{0}\approx 6$ TeV. The quartic coupling, $\lambda$, will be determined at loop level and we find remarkable agreement with the SM as discussed below.
The degree of fine-tuning of the theory is also suppressed by $\phi(0)$ in a subtle way. Rather than the naive result one would expect from the NJL model, $\delta g_{0}^{2}/g_{c}^{2}\sim|\mu|^{2}/M_{0}^{2}\sim 10^{-4}$, we now obtain a linear relation: $\delta g_{0}^{2}/g_{c}^{2}\sim|\mu|/M_{0}\sim 1\%$. To obtain these results we need the precise relationship between $g_{Y}$ and $\phi(0)$ which follows from the induced Yukawa coupling of the bound state to free fermions.
@@ II.5 The Induced Bound State Yukawa Interaction
We see in eq.(II.2) that the Yukawa interaction of the bound state with the free scattering state fermions is now induced from the $\sqrt{N_{c}}$ term:
@@
We therefore see that the induced Yukawa coupling to the field $H(x)$ is:
$g_{Y}=g_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0).$ (62)
Here the behavior of $\phi(0)$ is a suppression of $g_{Y}$ that is a power-law, $\sim\sqrt{|\mu|/M_{0}}$ for small $\mu$, near the critical coupling. This power-law behavior is significantly different than the slow RG (logarithmc) evolution in the old NJL-based top condensation model, which is why we now obtain $M_{0}\sim 6$ TeV and reduced fine-tuning $\sim|\phi(0)|^{2}\sim$ few $\%$ (rather than $M_{0}\sim 10^{15}$ GeV and extreme fine-tuning $\sim 10^{-26}$).
@@ III Quantum Fermion Loop Corrections to the Semiclassical Theory
$\!\!\!\!\!\!\!\!\!\!S_{eff}=\!\int\!d^{4}x\biggl{(}\widetilde{Z}DH^{\dagger}DH+M^{2}H^{\dagger}H-\frac{\lambda}{2}(H^{\dagger}H)^{2}\biggr{)},$ (64)
@@
Hence, in the present analysis we will use the NJL limit in obtaining approximate bilocal loop integral results, but judiciously replacing $\delta^{4}(x-y)\rightarrow M_{0}^{2}D_{F}(x-y)$. Physically, the bilocal theory has a true momentum space cutoff of loop integrals $\sim M_{0}^{2}$, due to $D_{F}(q^{2})$, while in the NJL theory this cut-off is implemented in loop integrals “by hand.” The replacement of the $\delta^{4}(x-y)$ by $M_{0}^{2}D_{F}(x-y)$ approximately implements the true physics of the bilocal theory, allowing us to obtain a potential that, in turn, allows the “back-reaction” of the solution. We are invoking the “indistinguishability” of the $\delta^{4}(x-y)$ from the bilocal $M_{0}^{2}D_{F}(x-y)$, but maintaining the bilocality of the present theory. By extending the auxiliary field equation to include the additional terms of $W(\Omega)$, then integrating out $B_{i}$, we obtain the quantum loop corrections to the semiclassical action. These are the analogues of the terms in eqs.(64,65) generated by eq.(63), but now for the bilocal fields.
@@
$\int_{xy}\biggl{(}g_{0}\overline{\omega}_{R}(y)D_{F}(x-y)\omega_{iL}(x)B^{i\dagger}(x,y)+h.c.\biggr{)}.$ (66)
@@
$+\!\int_{x...z^{\prime}}\!\!\!\![B^{\dagger}(x,y)B(x^{\prime},y^{\prime})]{\cal{G}}(x,y,x^{\prime},y^{\prime},w,z,w^{\prime},z^{\prime})[B^{\dagger}(w,z)B(w^{\prime},z^{\prime})]+$... (67)
@@ III.1 Main Loop
$-\frac{1}{2}g_{0}^{2}N_{c}\!\int_{x...y^{\prime}}\!\!\![B^{\dagger}(x,y)D_{F}(x-y)\mathop{\rm Tr}(S_{F}(x-x^{\prime})S_{F}(y^{\prime}-y))D_{F}(x^{\prime}-y^{\prime})B(x^{\prime},y^{\prime})],$ (68)
@@
$S_{F}(x)=\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{i\ell\!\!\!/\!\,\,}{\ell^{2}+i\epsilon}e^{i\ell\cdot x}\qquad\makebox{and}\qquad\mathop{\rm Tr}\biggl{(}\frac{(1-\gamma^{5})}{2}\ell\!\!\!/\!\,\,\;\ell\!\!\!/\!\,\,^{\prime}\biggr{)}=2\ell\cdot\ell^{\prime}.$ (69)
@@
$D_{F}(x-y)=-\int\frac{d^{4}q}{(2\pi)^{4}}\frac{e^{iq(x-y)}}{q^{2}-M_{0}^{2}}\rightarrow\int\frac{d^{4}q}{(2\pi)^{4}}\frac{e^{iq(x-y)}}{M_{0}^{2}}=\frac{1}{M_{0}^{2}}\delta^{4}(x-y),$ (70)
@@
Hence we have, with $\delta^{4}(x-y)=J\delta^{4}(2r)$,
@@
$\!\!\!\!\!\!\!\!\!\!=-\frac{g_{0}^{2}N_{c}}{M_{0}^{4}}\!\int_{XX^{\prime}}\![\widetilde{B}^{\dagger}(X,0)\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{d^{4}\ell^{\prime}}{(2\pi)^{4}}\frac{2\ell\cdot\ell^{\prime}}{\ell^{2}\ell^{\prime 2}}e^{i(\ell-\ell^{\prime})(X-X^{\prime})}\widetilde{B}(X^{\prime},0)].$ (74)
@@
$\qquad\qquad=\frac{ig_{0}^{2}N_{c}}{8\pi^{2}M_{0}^{4}}\!\int_{\overline{X}}[\widetilde{B}_{0}^{\dagger}\widetilde{B}_{0}]\biggl{(}M_{0}^{2}+P^{2}\ln\left(\frac{M_{0}}{m}\right)\biggr{)}.$ (75)
@@
We have obtained the analogues of the $M^{2}$ and $P^{2}$ terms in eqs.(64,65) acting on the auxiliary field. The $P^{2}$ term corresponds to a loop-generated kinetic term for the auxiliary field $B_{i}$, (see eq.(94)). In the framework of the NJL model and original top condensation theory, BHL , this corresponds the formation of a physical state, i.e., a resonance. We are interested presently in small $P^{2}$ so we will treat the $P^{2}\approx 0$ limit of the loop and return to this issue below.
@@
$W_{\cal{F}}(\Omega)=\frac{g_{0}^{2}N_{c}}{8\pi^{2}M_{0}^{2}}\!\int d^{4}X\;[\widetilde{B}^{\dagger}(X,0)\widetilde{B}(X,0)]\rightarrow\frac{g_{0}^{2}N_{c}J}{8\pi^{2}M_{0}^{2}}\!\int d^{4}Xd^{4}r\;\widetilde{B}^{\dagger}(X,r)\delta^{4}(2r)\widetilde{B}(X,r).$ (76)
@@
$W_{\cal{F}}(\Omega)=\!\int_{x...y^{\prime}}\!\!\!B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})\approx\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\int d^{4}x\;d^{4}y\;B^{\dagger}(x,y)D_{F}(x-y)B(x,y),$ (77)